OpenStudy (anonymous):

A 42 ball is fired horizontally with initial speed toward a 100 ball that is hanging motionless from a 1.1 -long string. The balls undergo a head-on, perfectly elastic collision, after which the 100 ball swings out to a maximum angle = 52.

OpenStudy (anonymous):

what is the question

OpenStudy (anonymous):

\[KE_0+PE_0=KE_f+PE_f\] \[h = l-(lcos(52)) \] \[(1/2)m_bv_0^2+0 = mgh \rightarrow mg(l-lcos(52))+(1/2)m_bv_f^2\] find vf or \[KE_0+KE_{angular_0} = KE_f+KE_{angilar_{final}}\]

OpenStudy (anonymous):

\[KE_{angular} = (1/2)I\omega^2\] \[I_{sphere} = (2/5)mr^2\]